Last week the price of Costco's (Ticker: COST) common stock closed at $302.50 per share. Use the data below to compute Costco's beta, then use the beta to estimate its required cost of equity. The prices have been adjusted for dividends and stock splits. Assume the appropriate risk-free rate is 3.50% and the market risk premium is 5.5%. Show your beta to 4 decimal places. Notice the ordering of the dates when estimating the beta.

Costco's Beta = ________      Costco's Cost of Equity Capital = _______ %

Friday Closing Prices for Costco (COST) and the S&P 500 Index

Week of

Costco

Friday Close

SP500 Friday Close

5/13/19

247.75

2859.53

5/20/19

246.75

2826.06

5/27/19

239

2752.06

6/3/19

255.5

2873.34

6/10/19

259.5

2886.98

6/17/19

265.5

2950.46

6/24/19

263.75

2941.76

7/1/19

267.75

2990.41

7/8/19

278.75

3013.77

7/15/19

281

2976.61

7/22/19

280.75

3025.86

7/29/19

272.5

2932.05

8/5/19

273

2918.65

8/12/19

273.5

2888.68

8/19/19

273.5

2847.11

8/26/19

294

2926.46

9/2/19

303.75

2978.71

9/9/19

292

3007.39

9/16/19

286.25

2992.07

9/23/19

286

2961.79

9/30/19

291.75

2952.01

10/7/19

297.5

2970.27

10/14/19

302.5

2997.95
Answer #1

Beta is given by the formula

B = Covariance CostCoReturns, SP500 Returns)/Variance(SP500 Returns)

F2 - X ✓ fc =COVARIANCE.P(D2:D23,E2:E23)/VAR.P(E2:E23) Beta of Costco =COVARIANCE.PD2:D23, E2:E23)/VAR.P(E2:E23) 7 1 Costco F

G2 - x fc =COVARIANCE.P(E2:E23,F2:F23)/VAR.P(F2:F23) в G 1 Week of Costco Friday Close SP500 Friday Close Costco Returns S&P

Cost of Equity is given by

Ke = R + B* (Rm-R)

where

Ke = Costo fequity

R= Risk Free Rate

B = Betaofthecompany

Rm = Market Returns

(Rm - Rp) = Market Risk Premium

=C2+C1*C3 1 Beta 2 Risk Free Rate 3 Market Risk Premium 4 Cost of Equity 1.2 10.035 0.055 =C2+C1*C3

(4 x ✓ с fix в 1 Beta 2 Risk Free Rate 3 Market Risk Premiun 4 Cost of Equity 1.2 3.50% 5.50% 10.100%